# Understanding the chemistry of temperate exoplanets atmospheres through experimental and numerical simulations

### v1

@@ 2.3 Mass Spectrometry
Mass Spectrometry is used as an analytical method to determine the molecular species produced. PAMPRE is equipped with a Hiden EQP system, which is used in residual gas analysis (RGA) mode, since only neutral species are considered in this study. A collector head is used to sample gaseous species through a 100 µm diameter orifice. The pressure inside the mass spectrometer is kept below 5x10-6 mbar. To study neutral species, an internal ionizer is used, with electron energy set at 70 eV. The resulting ions are then filtered and focused before entering the triple-filter mass analyzer, which detects masses from 1 to 200 m/z (mass-to-charge ratio) with a resolution of 1 atomic mass unit (amu). Finally, the ions arrive at an ion-counting detector — a secondary electron multiplier (SEM).
@@
Several molecular fragments at different m/z values are generated and form a ”fingerprint”, specific to each molecule and incident electron energy, known as the fragmentation pattern. A full mass spectrum thus corresponds to the superposition of the fragmentation patterns of all the species present in the analyzed gas mixture. Reference fragmentation patterns were measured directly in the lab for the reactive compounds, i.e. H2, CH4, CO and CO2. The reference fragmentation patterns of the products are extracted from the NIST Chemistry WebBook [^1] https://webbook.nist.gov/chemistry/ (NIST).
@@
All m/z over a selected mass range are scanned in BAR mode (Batched Analysis of Repetitions). This produces a histogram of peak intensities. Each mass spectrum contains 20 accumulated scans, over a mass range from 1 to 100 amu. A mass spectrum is taken at the initial state (spectrum OFF), when the gaseous mixture is injected inside the reactive chamber at 0.98 mbar, while the plasma discharge is still switched off. The plasma is switched on and after five minutes, a spectrum is taken (spectrum ON). To ensure that we have reached a stable steady state, we perform multiple ion detection (MID) when we activate the plasma. This allows us to track several m/z ratios over time, thereby monitoring consumption/production and determining when we reach a stable state. In all our experiments, we observe stabilization after a maximum of 200 seconds.
@@
We can see the evolution of the gas composition by plotting the superposition of the two spectra ON and OFF (Figure 2), or plotting the difference of intensity between these two spectra (ON-OFF) to see the production and destruction of chemical species.
@@
Figure 2: Mass spectrum of a gas mixture of 95% H2 + 5% CH4 before (gold) and during irradiation (blue).
@@
The measured mass spectra are processed first by removing the noise from the data: all intensities below 10 counts/s are removed, as the instrument’s sensitivity is 106 counts/s maximum. Each spectrum is then normalized by its maximum intensity. In the direct-flow experiments, the maximum intensity is always at m/z 2, corresponding to H2. In trap-and-release experiments, the maximum intensity depends on the composition. Mean values and standard deviation at 95% are calculated from the 20 cycles for each m/z. If the difference between the average intensity and the standard deviation is below 10-6, it is considered as noise and removed from the data.
@@
While the reconstruction of a mass spectrum with fragmentation patterns works well with few compounds, it becomes much more difficult to analyze and identify species in more complex mixtures with dozens of products. This is due to the superposition of several fragmentation patterns and the variability associated with these patterns — which depends on the geometry of the ion source, the mass-dependent filtering and detection efficiency of the instrument. We therefore need a complementary analytical method in order to make unambiguous identifications.
@@ 2.4 Infrared spectroscopy
IR spectroscopy is used to help identify species in CH4/CO/CO2 mixtures. The PAMPRE setup is connected to a multipass infrared white cell with an optical beam length of 10 m. The cell is positioned in the sample compartment of an FTIR spectrometer (Nicolet 6700, Thermo Fisher).
Gas mixtures flow directly through this cell to characterize the dissociation rate of reactants and the formation of major products. However, the concentration of minor products is too low relative to the instrument’s thresholds to allow their identification. To overcome this limit, we use a cold trapping strategy to concentrate our products before IR identification. A glass trap installed at the outlet of the system is immersed in liquid nitrogen (78 K) (see Figure 1), ensuring the entire flow passes through it, condensing most of the gases. H2, CH4 and CO, with condensation temperatures below 78 K, are not trapped, which helps to concentrate minor products. However, precise quantification is not possible due to this bias-inducing differentiated condensation, but robust product identification is already an important step forward. The irradiated gas stream is allowed to pass through the trap for 20 min; trapping is then stopped, the trap is heated to sublimate the condensed species, the gases are released into the multipass cell, and heated to 70°C so that water vapor does not condense on the mirrors. The pressure is measured inside the cell and it does not exceed a few mbar. Since there is no longer any irradiation, we remain at low pressure and do not reach very high temperatures, so we expect the kinetics to be slow and no significant new reactions to occur. Infrared spectra are measured through the software OMNIC (Thermo Scientific), with a resolution of 0.5 cm-1using the Graymann-Harris apodisation function Harris_1978. 150 scans were accumulated for each measurement, averaged, and combined into a spectrum.
Preprocessing the spectra, using OMNIC software, transforms the single-beam data acquired into transmittance or absorbance spectra, by dividing them by blank spectra previously acquired under high vacuum conditions (less than 10-5 mbar) and by applying a manual baseline correction. The IR spectra are compared with those calculated at the experimental pressure (1 mbar) and temperature (340 K), based on the opacities of several species (Figure 12). The opacities are taken from the databases DACE  [^2] https://dace.unige.ch/opacityDatabase/? and HITRAN (Gordon2017_HITRAN). However, these identifications are limited by the availability of molecules — organic compounds with 3 or more carbon atoms are not included. For more complex molecules, identification is performed using the NIST database, but this does not allow for temperature and pressure dependence to be taken into account.
@@ 2.5 Atmospheric numerical simulations
To model the chemistry occurring in the reactor chamber, we used the 0D photochemical model ReactorUI (Peng_2014; Pernot_2023), as described in Peng_2014. This complementary approach to experimental characterization helps to identify the species produced and to elucidate the main chemical pathways leading to their formation (Bourgalais_2020).
@@
The chemical network was originally developed for the CHN atmosphere of Titan with neutral-neutral reactions (Hebrard_2007; Hebrard_2009; Dobrijevic_2016), ions (Carrasco_2008; Plessis_2012) and photo-processes (from the Leiden (Heays_2017) and SWRI (Huebner_2015) databases). For the studies of the formation of more complex oxygenated molecules, such as H2CO and CH3OH, we extended the chemical network with that of CHO VULCAN (Tsai_2021), specific to sub-Neptunes, based on Venot_2020a.
Simulations of each experimental composition were run. Considering chemical rate uncertainties with a Monte Carlo approach, we perform 500 runs for each configuration. This Monte Carlo approach to characterizing uncertainties has already been used in several atmospheric models of solar system objects, such as Triton (250 runs, Benne_2022), Titan (500 runs, Hebrard_2009) or Neptune (1000 runs, Dobrijevic_2010). As a result, we evaluate the final species abundance values with its uncertainty. These simulations also provide information on the contributions of specific reactions to the formation or consumption of a given molecule, enabling reaction paths to be extracted.
@@ 3.1.1 Reduced organic compounds
We observe the formation of reduced organic compounds, composed only of carbon and hydrogen atoms, forming chains with up to 6 carbon atoms. In mass spectrometry, groups of peaks form at m/z up to 80 when a 95% H2 + 5% CH4 gas mixture is irradiated (Figure 2). The specific fragmentation patterns (FP) of reduced compounds with 2 and 3-carbon atoms (C2 and C3) are shown in Figure 11. These patterns are normalized to the mass spectrum measured in the irradiated gas mixture of 95% H2 + 5% CH4, and directly compared with the corresponding mass spectrum. In the case of reduced atmospheres, C2 identification using MS alone is possible, particularly for C2H2 and C2H6, which exhibit strong characteristic peaks. Infrared spectroscopy is used to determine the nature of the chemical bonds between carbon atoms, distinguishing between single, double and triple bonds. The IR features used to identify acetylene (C2H2), ethene (C2H4), propyne (C3H4) and propene (C3H6) are shown in Figure 3.
@@ 3.1.2 Oxidized organic compounds
Using IR, we can see in CO and CO2-rich mixtures the robust formation of the oxidized organic compounds methanol (CH3OH), formaldehyde (H2CO) and acetaldehyde (CH3CHO). The features used for their identification are shown in Figure 4, at 1033 cm-1(9.68 µm) for CH3OH and at 1745 cm-1(5.73 µm) for the aldehydes (H2CO and CH3CHO). In MS, the production of CH3OH is identified by the appearance of the characteristic peaks at m/z 31 and 32. The fragmentation patterns of H2CO and CH3CHO overlap with those of C2H6 and C3H6, which makes their identification only through MS a source of uncertainty.
@@ 3.1.3 Other oxidized compounds
(a) IR feature at 730 cm-1(13.70 µm), attributed to C2H2.
(b) IR feature at 950 cm-1(10.53 µm), attributed to C2H4.
(c) IR feature at 2150 cm-1(4.65 µm), attributed to C3H4.
(d) IR feature at 912 cm-1(10.96 µm), attributed to C3H6.
@@
(a) IR feature at 1033 cm-1(9.68 µm), attributed to CH3OH.
(b) IR feature at 1745 cm-1(5.73 µm), attributed to aldehydes.
@@ 3.2.1 Reduced organic compounds
Figure 6: Predicted abundance values with 0D-simulations of the 2-carbon atoms reduced organics in the different gas mixtures.
The formation of reduced organics is observed under all experimental conditions, with the appearance of carbon chain features both in MS and IR as presented in section 3.1, shown in Figures 2 and 3, and detailed in Table 2.
The organic growth is most significant in reduced mixtures, and increases with methane concentration, and hence with metallicity, as expected and already highlighted by previous experimental studies (Horst_2018). The mass spectra show an increase in the intensity of all peaks with increasing methane content, as shown in Figure 5(a). We observe peak groups characteristic of compounds comprising up to 4-carbon atoms (C4), at m/z 50-60. Although it is difficult to quantify the species precisely, the relative intensities of the peaks relating to C2 compounds appear to be of comparable magnitudes, reflecting similar concentrations. Based on the 0D simulations performed with ReactorUI, we observe that the abundance of C2H2 is approximately two orders of magnitude lower than that of C2H4 and C2H6, and decreases with increasing CH4 concentration (Figure 6(a)). In contrast, the abundance of C2H4 and C2H6 increase with increasing methane concentration.
The formation of reduced organics also occurs in oxidized mixtures, although it is more limited than in reduced mixtures, as previously highlighted in particular for CO2-rich mixtures (Trainer_2004). It is observed in MS, as we find lower intensities of C2 (m/z 20-30) and C3 (m/z 30-40) specific peak groups in CO and CO2-rich mixtures (Figures 5(b) and 5(d)) than in reduced mixtures (Figure 5(a)). In Figures 5(b) and 5(d), we do not observe the C4 specific peak groups at m/z 50-60 that were present in Figure 5(a), reflecting inhibited organic growth. No selectivity is observed in oxidized mixtures between the C2 compounds, with simulated abundance values similar for C2H2, C2H4 and C2H6 (Figure 6(a),6(b)).
The organic growth is predicted by the 0D-simulations to be more efficient with CO-rich mixtures than with CO2-rich mixtures, as shown in Figure 6(a) where we see that the abundance values of 2-carbon chains are higher with CO than with CO2. This is not obvious with the MS results presented in Figures 5(b) and 5(d), where we see similar relative intensities of the characteristic m/z ratios of compounds C2 and C3. However this is supported by the IR data, as we observe the appearance of species absorption peaks of 3-carbon chains only in CO-rich mixtures (CO and CH4/CO mixtures). For example, we only see the 912 cm-1(10.96 µm) peak attributed to C3H6 clearly appearing in CH4/CO mixtures, as shown in Figure 3(d) (solid lines).
@@
When H2, CH4 and an oxidized compound —CO or CO2 — are merged, the efficiency of organic growth is a hybrid of the two previous cases. It is increased compared to pure oxidized mixtures, and slightly decreased compared to pure reduced mixtures. In the mass spectra, characteristic peak groups of compounds C2 and C3 are seen in Figures 5(c) and 5(e), at higher intensities than in the oxidized mixtures presented above, but lower than in the reduced mixtures. Figure 6(b) shows the effect of adding methane to the atmospheres, particularly for the formation of C2H4 and C2H6: the gain is 2 to 6 orders of magnitude, selectively favoring the formation of C2H4 and C2H6 over C2H2. The more oxidized the composite atmosphere is, the less efficient the formation of reduced organics is; this is visible in Figure 6(c), where the predicted abundance values for the 94% H2 + 5% [ CO or CO2 ] + 1% CH4 mixtures are the lowest. We also observe in Figure 6(d) that when 0.5% CO2 is added to CH4/CO mixtures, the reduced organics abundance values decrease.

### v2

@@ 2.3 Mass spectrometry
Mass spectrometry is used as an analytical method to determine the molecular species produced. PAMPRE is equipped with a Hiden EQP system, which is used in residual gas analysis (RGA) mode, since only neutral species are considered in this study. A collector head is used to sample gaseous species through a 100 µm diameter orifice. The pressure inside the mass spectrometer is kept below 5x10-6 mbar. To study neutral species, an internal ionizer is used, with electron energy set at 70 eV. The resulting ions are then filtered and focused before entering the triple-filter mass analyzer, which detects masses from 1 to 200 m/z (mass-to-charge ratio) with a resolution of 1 atomic mass unit (amu). Finally, the ions arrive at an ion-counting detector — a secondary electron multiplier.
@@
Several molecular fragments at different m/z values are generated and form a ”fingerprint” specific to each molecule and incident electron energy, known as the fragmentation pattern. A full mass spectrum thus corresponds to the superposition of the fragmentation patterns of all the species present in the analyzed gas mixture. Reference fragmentation patterns were measured directly in the lab for the reactive compounds, i.e., H2, CH4, CO and CO2. The reference fragmentation patterns of the products are extracted from the NIST Chemistry WebBook [^1] https://webbook.nist.gov/chemistry/ (Linstrom and Mallard, 2001).
@@
All m/z over a selected mass range are scanned in BAR mode (Batched Analysis of Repetitions). This produces a histogram of peak intensities. Each mass spectrum contains 20 accumulated scans, over a mass range from 1 to 100 amu. A mass spectrum is taken at the initial state (spectrum OFF), when the gaseous mixture is injected inside the reactive chamber at 0.98 mbar, while the plasma discharge is still switched off. The plasma is switched on and after five minutes, a spectrum is taken (spectrum ON). To ensure that we had reached a stable steady state, we performed multiple ion detection when we activated the plasma. This allowed us to track several m/z ratios over time, thereby monitoring consumption and production and determining when we reach a stable state. In all our experiments, we observed stabilization after a maximum of 200 seconds.
@@
The evolution of the gas composition can be seen by our plotting of the superposition of the two ON and OFF spectra (Figure 2). Our plotting of the difference of intensity between these two spectra (ON-OFF) reveals the production and destruction of the chemical species.
@@
Figure 2: Mass spectrum of a gas mixture of 95% H2 plus 5% CH4 before (gold) and during irradiation (blue).
@@
The measured mass spectra are processed first by removing the noise from the data: all intensities below 10 counts/s are removed, as the instrument’s sensitivity is 106 counts/s maximum. Each spectrum is then normalized by its maximum intensity. In the direct-flow experiments, the maximum intensity is always at m/z 2 amu, corresponding to H2. In trap-and-release experiments, the maximum intensity depends on the composition. Mean values and standard deviation at 95% are calculated from the 20 cycles for each m/z. If the difference between the average intensity and the standard deviation is below 10-6, it is considered as noise and removed from the data.
@@
While the reconstruction of a mass spectrum with fragmentation patterns works well with a few compounds, it becomes much more difficult to analyze and identify species in more complex mixtures with dozens of products. This is due to the superposition of several fragmentation patterns and the variability associated with these patterns — which depends on the geometry of the ion source, the mass-dependent filtering, and detection efficiency of the instrument. We therefore need a complementary analytical method in order to make unambiguous identifications.
@@ 2.4 Infrared spectroscopy
Infrared spectroscopy was used to help identify species in CH4/CO/CO2 mixtures. The PAMPRE setup was connected to a multipass infrared white cell with an optical beam length of 10 m. The cell was positioned in the sample compartment of an FTIR spectrometer (Nicolet 6700, Thermo Fisher).
Gas mixtures flow directly through this cell to characterize the dissociation rate of reactants and the formation of major products. However, the concentration of minor products is too low relative to the instrument’s thresholds to allow their identification. To overcome this limit, we use a cold trapping strategy to concentrate our products before IR identification. A glass trap installed at the outlet of the system is immersed in liquid nitrogen (78 K) (see Figure 1), ensuring the entire flow passes through it, condensing most of the gases. H2, CH4 and CO, with condensation temperatures below 78 K, are not trapped, which helps to concentrate minor products. However, precise quantification is not possible due to this bias-inducing differentiated condensation, but robust product identification is already an important step forward. The irradiated gas stream is allowed to pass through the trap for 20 min; trapping is then stopped, the trap is heated to sublimate the condensed species, the gases are released into the multipass cell, and heated to 70°C so that water vapor does not condense on the mirrors. The pressure is measured inside the cell and it does not exceed a few millibars. Since there is no longer any irradiation, we remain at low pressure and do not reach very high temperatures, so we expect the kinetics to be slow and no significant new reactions to occur. Infrared spectra are measured through the software OMNIC (Thermo Scientific), with a resolution of 0.5 cm-1using the Graymann-Harris apodization function Harris (1978). 150 scans were accumulated for each measurement, averaged, and combined into a spectrum.
Preprocessing the spectra, using OMNIC software, transforms the single-beam data acquired into transmittance or absorbance spectra, by dividing them by blank spectra previously acquired under high vacuum conditions (less than 10-5 mbar) and by applying a manual baseline correction. The IR spectra were compared with those calculated at the experimental pressure (1 mbar) and temperature (340 K), based on the opacities of several species (Figure 12). The opacities are taken from the databases DACE  [^2] https://dace.unige.ch/opacityDatabase/? and HITRAN (Gordon et al., 2017). However, these identifications were limited by the availability of molecules — organic compounds with three or more carbon atoms are not included. For more complex molecules, identification is performed using the NIST database, but this does not allow for temperature and pressure dependence to be taken into account.
@@ 2.5 Atmospheric numerical simulations
To model the chemistry occurring in the reactor chamber, we used the 0D photochemical model ReactorUI (Peng et al., 2014; Pernot, 2023), as described in Peng et al. (2014). This complementary approach to experimental characterization helps to identify the species produced and to elucidate the main chemical pathways leading to their formation (Bourgalais et al., 2020).
@@
The chemical network was originally developed for the CHN atmosphere of Titan with neutral-neutral reactions (Hébrard et al., 2007, 2009; Dobrijevic et al., 2016), ions (Carrasco et al., 2008; Plessis et al., 2012) and photo-processes (from the Leiden (Heays et al., 2017) and SWRI (Huebner and Mukherjee, 2015) databases). For the studies of the formation of more complex oxygenated molecules, such as H2CO and CH3OH, we extended the chemical network with that of CHO VULCAN (Tsai et al., 2021), specific to sub-Neptunes, based on Venot et al. (2020).
Simulations of each experimental composition were run. Considering chemical rate uncertainties with a Monte Carlo approach, we perform 500 runs for each configuration. This Monte Carlo approach to characterizing uncertainties has already been used in several atmospheric models of Solar System objects, such as Triton (250 runs, Benne et al. (2022)), Titan (500 runs, Hébrard et al. (2009)) or Neptune (1000 runs, Dobrijevic et al. (2010)). As a result, we evaluate the final species abundance values with its uncertainty. These simulations also provide information on the contributions of specific reactions to the formation or consumption of a given molecule, enabling reaction paths to be extracted.
@@ 3.1.1 Reduced organic compounds
We observe the formation of reduced organic compounds, composed only of carbon and hydrogen atoms, forming chains with up to six carbon atoms. In MS, groups of peaks form at m/z up to 80 when a 95% H2 + 5% CH4 gas mixture is irradiated (Figure 2). The specific fragmentation patterns (FP) of reduced compounds with two and three carbon atoms (C2 and C3) are shown in Figure 11. These patterns are normalized to the mass spectrum measured in the irradiated gas mixture of 95% H2 + 5% CH4, and directly compared with the corresponding mass spectrum. In the case of reduced atmospheres, C2 identification using MS alone is possible, particularly for C2H2 and C2H6, which exhibit strong characteristic peaks. Infrared spectroscopy is used to determine the nature of the chemical bonds between carbon atoms, distinguishing between single, double and triple bonds. The IR features used to identify acetylene (C2H2), ethene (C2H4), propyne (C3H4) and propene (C3H6) are shown in Figure 3.
@@ 3.1.2 Oxidized organic compounds
Using IR, we could see in CO and CO2-rich mixtures the robust formation of the oxidized organic compounds methanol (CH3OH), formaldehyde (H2CO) and acetaldehyde (CH3CHO). The features used for their identification are shown in Figure 4, at 1033 cm-1(9.68 µm) for CH3OH and at 1745 cm-1(5.73 µm) for the aldehydes (H2CO and CH3CHO). In MS, the production of CH3OH is identified by the appearance of the characteristic peaks at m/z 31 and 32. The fragmentation patterns of H2CO and CH3CHO overlap with those of C2H6 and C3H6, which makes their identification only through MS a source of uncertainty.
@@ 3.1.3 Other oxidized compounds
(a) Infrared feature at 730 cm-1(13.70 µm), attributed to C2H2.
(b) Infrared feature at 950 cm-1(10.53 µm), attributed to C2H4.
(c) Infrared feature at 2150 cm-1(4.65 µm), attributed to C3H4.
(d) Infrared feature at 912 cm-1(10.96 µm), attributed to C3H6.
@@
(a) Infrared feature at 1033 cm-1(9.68 µm), attributed to CH3OH.
(b) Infrared feature at 1745 cm-1(5.73 µm), attributed to aldehydes.
@@ 3.2.1 Reduced organic compounds
Figure 6: Predicted abundance values with 0D simulations of the two-carbon atoms reduced organics in the different gas mixtures.
The formation of reduced organics is observed under all experimental conditions. We noted the appearance of carbon chain features both in MS and IR as presented in section 3.1, shown in Figures 2 and 3, and detailed in Table 2.
The organic growth is most significant in reduced mixtures, and it increases with methane concentration and hence with metallicity, as expected and already highlighted by previous experimental studies (Hörst et al., 2018). The mass spectra show an increase in the intensity of all peaks with increasing methane content, as shown in Figure 5(a). We observed peak groups characteristic of compounds comprising up to four carbon atoms (C4), at m/z 50-60. Although it is difficult to quantify the species precisely, the relative intensities of the peaks relating to C2 compounds appear to be of comparable magnitudes, reflecting similar concentrations. Based on the 0D simulations performed with ReactorUI, we observe that the abundance of C2H2 is approximately two orders of magnitude lower than that of C2H4 and C2H6, and decreases with increasing CH4 concentration (Figure 6(a)). In contrast, the abundance of C2H4 and C2H6 increase with increasing methane concentration.
The formation of reduced organics also occurs in oxidized mixtures, although it is more limited than in reduced mixtures, as previously highlighted in particular for CO2-rich mixtures (Trainer et al., 2004). It is observed in MS, as we find lower intensities of C2 (m/z 20-30) and C3 (m/z 30-40) specific peak groups in CO and CO2-rich mixtures (Figures 5(b) and 5(d)) than in reduced mixtures (Figure 5(a)). In Figures 5(b) and 5(d), we do not observe the C4 specific peak groups at m/z 50-60 that were present in Figure 5(a), reflecting inhibited organic growth. No selectivity is observed in oxidized mixtures between the C2 compounds, with simulated abundance values similar for C2H2, C2H4 and C2H6 (Figure 6(a),6(b)).
The organic growth is predicted by the 0D-simulations to be more efficient with CO-rich mixtures than with CO2-rich mixtures, as shown in Figure 6(a) where we see that the abundance values of two-carbon chains are higher with CO than with CO2. This is not obvious with the MS results presented in Figures 5(b) and 5(d), where we see similar relative intensities of the characteristic m/z ratios of compounds C2 and C3. However this is supported by the IR data, as we observe the appearance of species absorption peaks of three-carbon chains only in CO-rich mixtures (CO and CH4/CO mixtures). For example, we only see the 912 cm-1(10.96 µm) peak attributed to C3H6 clearly appearing in CH4/CO mixtures, as shown in Figure 3(d) (solid lines).
@@
When H2, CH4 and an oxidized compound — CO or CO2 — are merged, the efficiency of organic growth is a hybrid of the two previous cases. It is increased compared to pure oxidized mixtures, and slightly decreased compared to pure reduced mixtures. In the mass spectra, characteristic peak groups of compounds C2 and C3 are seen in Figures 5(c) and 5(e), at higher intensities than in the oxidized mixtures presented above, but lower than in the reduced mixtures. Figure 6(b) shows the effect of adding methane to the atmospheres, particularly for the formation of C2H4 and C2H6: the gain is two to six orders of magnitude, selectively favoring the formation of C2H4 and C2H6 over C2H2. The more oxidized the composite atmosphere is, the less efficient the formation of reduced organics is. This relationship is visible in Figure 6(c), where the predicted abundance values for the 94% H2 + 5% [ CO or CO2 ] + 1% CH4 mixtures are the lowest. We also observe in Figure 6(d) that when 0.5% CO2 is added to CH4/CO mixtures, the reduced organics abundance values decrease.